OpenStudy (anonymous):

which ratio forms a proportion with 30/42 15/27 28/40 45/63 21/15

OpenStudy (jusaquikie):

\[\frac{ 30 }{42 } (\div) 6 = \frac{ 5 }{7 }\] \[\frac{ 45 }{63 }(\div)9 = \frac{ 5 }{7 }\] so the answer is \[\frac{ 45 }{63}\]

OpenStudy (jusaquikie):

another way to get it is to cross multiply if you get the same number they are proportional. 30*63 = 1890 45*42 = 1890

OpenStudy (anonymous):

thanks!

OpenStudy (jusaquikie):

np good luck

OpenStudy (anonymous):

much needed ha.